OpenStudy (anonymous):

Please help. -Remember compounded twice a year. :/ @jim_thompson5910

jimthompson5910 (jim_thompson5910):

use the formula A = P*(1+r/n)^(n*t) in this case, P = 4280.83 r = 0.054 n = 2 t = 25 plug them into the formula to find the value of A. Tell me what you get.

OpenStudy (anonymous):

A= 16219. 97

jimthompson5910 (jim_thompson5910):

That's the total amount of money in the account after 25 years. That money consists of the initial principal ($4,280.83) plus interest earned over the 25 years. Let x be the amount of interest earned. So that means P + I = A 4280.83+x = 16219. 97 x = ??

OpenStudy (anonymous):

x= 11939.14

jimthompson5910 (jim_thompson5910):

that's your answer

jimthompson5910 (jim_thompson5910):

that's the amount of interest (and just interest alone) you've earned

OpenStudy (anonymous):

My final answer? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes that's what they want

OpenStudy (anonymous):

Thank you!